# The number of cultural traits, evolving genealogies, and the descendant process

### v1

@@ 3.3.3 Evolving genealogies
Indeed, the realisation shows the sawtooth picture with linear increase of $(L_{N}^{t})_{t\geq 0}$ interrupted by sudden near-extinctions at random times. But to understand it, we need the finer picture of $\lvert\Lambda^{t}_{[n]}(\tau)\rvert_{\tau\geq 0}$ in Figure 14. The figure shows realisations of $\lvert\Lambda^{t_{i}}_{[n]}(t_{i}-t)\rvert_{t\leq t_{i}}$ for $i\in\{1,2,3,4\}$, $n\in\{1,N\}$, and starting times $t_{i}$ chosen so that
@@
where $t_{i}^{\rm e}$ is the time point where $(\Lambda^{t_{i}}_{[N]}(t_{i}-t)_{t\leq t_{i}})$ goes extinct; so $t_{i}-t_{i}^{\rm e}$ is the corresponding realisation of $T^{t_{i}}$. Note that, due to (34), the $t_{i}-t_{i}^{\rm e}$ are coupled for $i\in\{2,3,4\}$. With $\lesssim$ and $\ll$, we indicate that the quantities are close and not close to each other, respectively. We now see that the realisations of $(\Lambda^{t_{3}}_{[N]}(t_{3}-t))_{t\leq t_{3}}$ and $(\Lambda^{t_{2}}_{[N]}(t_{2}-t))_{t\leq t_{2}}$ quickly get close to, and actually become identical with, the realisation of $(\Lambda^{t_{4}}_{[N]}(t_{4}-t))_{t\leq t_{4}}$; in particular, they join into the same metastable set and are extinguished with it. Likewise, the realisations of $(\Lambda^{t_{i}}_{\alpha}(t_{i}-t))_{t\leq t_{i}}$, $i\in\{2,3,4\}$ and $\alpha\in[N]$, either die out quickly or coalesce with the metastable set. This coalescence is clear by arguments similar to those in Section 3.3.2: since $\Lambda^{t_{4}}_{[N]}(t_{4}-t_{3})\subseteq\Lambda^{t_{3}}_{[N]}(0)=[N]$, we have $\Lambda^{t_{4}}_{[N]}(t_{4}-t_{3}+\tau)\subseteq\Lambda^{t_{3}}_{[N]}(\tau)$ for all $\tau>0$; so, by moving towards the metastable size, the latter process joins into the former from above. Likewise, $\Lambda^{t_{3}}_{\alpha}(0)\subseteq\Lambda^{t_{3}}_{[N]}(0)$, so $\Lambda^{t_{3}}_{\alpha}(\tau)\subseteq\Lambda^{t_{3}}_{[N]}(\tau)$ for $\tau>0$; hence the former process, if it does not die out quickly, joins into the latter from below. Analogous arguments hold for the other time points.
@@
So far, we have followed the ancestral process backward in time. But the figure can also be read forward. Let $\Gamma_{[n]}^{t}(\sigma)$ be the set of descendants at time $t+\sigma$, $\sigma\geq 0$, of the set $[n]$ of individuals at forward time $t$. The process $(\Gamma_{[n]}^{t}(\sigma))_{\sigma\geq 0}$ has the same law as $(\Lambda_{[n]}^{t}(\tau))_{\tau\geq 0}$ with $\tau$ replaced by $\sigma$; this follows immediately from the fact that the set of descendants (ancestors) is obtained by following all learning arrows in the forward (backward) direction (and pruning lines that meet a cross), and a reversal of all arrows does not change the law of the process. As a side remark, let us note that $\lvert\Gamma_{[n]}^{0}(t)\rvert_{t\geq 0}$ has the same law as $(X(t))_{t\geq 0}$ of our single-trait model with $X(0)=n$; this reflects the self duality of the SIS model. In any case, like $\lvert\Lambda_{[N]}^{t}(\tau)\rvert_{\tau\geq 0}$, also $\lvert\Gamma_{[N]}^{t}(\sigma)\rvert_{\sigma\geq 0}$ moves to the metastable state around $N\bar{\xi}$ quickly, and like $\lvert\Lambda_{\alpha}^{t}(\tau)\rvert_{\tau\geq 0}$, also $\lvert\Gamma_{\alpha}^{t}(\sigma)\rvert_{\sigma\geq 0}$ dies out quickly with probability $u/s$ and otherwise moves to the metastable state.
@@
[…]
@@
$\Lambda_{[N]}^{t}(t-s)=\{\alpha:\Gamma_{\alpha}^{s}(t-s)\neq\varnothing\}.$ (35)
@@
In general, $\Lambda_{[N]}^{t}(t-s)\neq\Gamma_{[N]}^{s}(t-s)$. However, it is true that
$\Lambda_{[N]}^{t}(t-s)=\varnothing\Longleftrightarrow\Gamma_{[N]}^{s}(t-s)=\varnothing$ (36)
@@
for any $0<s<t$, because if the population at time $t$ does not have ancestors at time $s$, then the descendants of the population at time $s$ do not survive until $t$. Moreover, if $t\gg s$ and the sets in (36) are not empty, we have
$\lvert\Lambda_{[N]}^{t}(t-s)\rvert\approx N\bar{\xi}\approx\lvert\Gamma_{[N]}^{s}(t-s)\rvert$ (37)
@@
due to metastability. In the realisation of the figure, the equalities in (36) are true for $s=t_{1}$, $t=t_{2}$; that is, both processes go extinct in $[t_{1},t_{2}]$. In contrast, for $s=t_{2}$ and $t=t_{3}$ or $t=t_{4}$, (37) applies.
Now, (36) and (37) together allow us to rewrite $L_{N}^{t}$ of (32) as
$L_{N}^{t}=\int_{0}^{t}\lvert\Gamma_{[N]}^{s}(t-s)\rvert\,\mathrm{d}s=\int_{0}^{S^{t}}\lvert\Gamma_{[N]}^{t-\sigma}(\sigma)\rvert\,\mathrm{d}\sigma\approx S^{t}N\bar{\xi},$ (38)
@@
where $S^{t}$ denotes the largest $\sigma\leq t$ such that $\Gamma_{[N]}^{t-\sigma}(\sigma)\neq\varnothing$, or, equivalently, the smallest $\sigma\leq t$ such that $\Gamma_{[N]}^{(t-\sigma)-}(\sigma)=\varnothing$. This is the time [^6] Note that $(\Gamma_{[n]}^{t}(\sigma))_{\sigma\geq 0}$ is based on the forward process, so it is càdlàg; that is, a jump at time $t$ means that the ’old’ state applies until time $t-$ (the moment ‘just before’ $t$), and at time $t$, the process is already in the ‘new’ state. This implies that the ancestors are extinct at $(t-S^{t})-$, but alive at $t-S^{t}$. since the last extinction event before $t$ (assumed to be $t$ if no extinction has taken place before $t$). In other words, the population at forward time $t$ has their oldest ancestor(s) at forward time $t-S^{t}$. Since the extinction times are distinct random points on the time axis, $(S^{t})_{t\geq 0}$ is a sawtooth function that increases linearly with slope 1 and is reset to some small value at random times. So (38) explains the linear increase in forward time, interrupted by steep descents at the jumps of $(S_{t})_{t\geq 0}$.
@@
We can now also understand the dynamics of $(L_{1}^{t})_{t\geq 0}$ as shown in the lower panel of Figure 13. Since, for any given $t$, $(\Lambda^{t}_{\alpha}(\tau))_{\tau\geq 0}$ quickly either dies out or gets close to $(\Lambda^{t}_{N}(\tau))_{\tau\geq 0}$, it is clear that, most of the time, $(L_{1}^{t})_{t\geq 0}$ is either close to 0 or close to $(L_{N}^{t})_{t\geq 0}$. The frequent, random transitions between the two possibilities come from the fact that, at the metastable state over time, the actual set of lines that has old ancestors, while having approximately constant size, moves around in the population via the events of teaching individuals outside the current set, and by death events, thus rapidly including or excluding individuals. That is, by learning from someone with old ancestors, an individual acquires the entire ancestry of the parent; and when the individual dies, it loses all its ancestry.
Let us finally get back to the traits via the relationship $C_{n}(t)\sim\text{Poi}(\mu L_{n}(t))$ and the observation that the variability in $C_{n}$ is largely governed by the variability of $L_{n}$ rather than that of the innovation process. It is then no surprise that the time series $(\mu\ell_{N}^{t})_{t\geq 0}$ and $(c_{N}(t))_{t\geq 0}$ agree almost perfectly in our simulations, and likewise for $(\mu\ell_{1}^{t})_{t\geq 0}$ and $(c_{1}(t))_{t\geq 0}$, as also illustrated in Figure 13. This confirms that the dynamics of the traits results from the genealogical features of the model. In particular, traits are collected at a constant rate while the genealogy is in a metastable state, and most traits are lost simultaneously when the genealogy collapses.
@@
Let us note in passing that, for the parameter values in Figure 13, (31) gives $\mathbb{E}\left[T\right]\approx 85.3$. Manual counting in Figure 13 yields $\approx 83$ sawteeth, which amounts to a mean tooth length (or extinction time) of 10000/83 $\approx$ 120. The overestimation of the length is presumably due to the underestimation of the number of teeth because small teeth are not resolved.
@@ 3.4 Popularity spectrum
Aoki (2018) (see also Strimling et al. (2009); Fogarty et al. (2015, 2017)) studied the popularity spectrum $({P_{l}})_{l\in[N]}$, where $P_{l}\mathrel{\mathop{:}}=\mathbb{E}\left[C_{N,l}\right]$ and $C_{n,l}$ is the number of traits carried by exactly $l$ individuals in a sample of size $n$ at stationarity. The aforementioned papers obtained the popularity spectrum in a discrete-time model via the equilibrium condition in the forward model. In Appendix D, we use the corresponding continuous-time version to derive
@@ 4 Discussion
Mathematical models for the accumulation of cultural traits in a population over time are crucial for understanding the cultural variation in human populations. Here we have investigated cultural dynamics in which selectively neutral, discrete cultural traits are invented by individuals independently of one another at a constant rate and transmitted between individuals through random social learning. Related previous analyses (Strimling et al., 2009; Lehmann et al., 2011; Fogarty et al., 2015, 2017; Aoki, 2018) were mostly static and centered around the traits, focussing on the expectation of quantities like the number of distinct cultural traits maintained in a population at stationarity. Our approach is different in three ways: first, we concentrate on the underlying genealogies (as first described by Aguilar and Ghirlanda (2015)) and thus obtain results that remain true independently of the traits; and second, we investigate the dynamical aspects, which seem to be unexplored so far. At the heart of this are the time evolution of the set of ancestors of a single individual or the entire population; the coalescent process between the various ancestral sets and the metastability of the corresponding counting processes, which agree with the well-studied stochastic logistic model (Nåsell, 2011; Foxall, 2021); and the concept of evolving genealogies. Third, we also obtain moments of $L_{n}$ and $C_{n}$.
@@
To keep the mean number of traits transmitted via learning events constant, we replace $s$ by $s/b$. For $b=0.99$ and, hence, our usual choice $s=1.3$ replaced by $s=130/99$, $(c_{N}(t))_{t\geq 0}$ still follows $\mu(\ell_{N}^{t})_{t\geq 0}$ closely, see Figure 16. For $b=13/14$ and, hence, $s=1.3$ replaced by $s=1.4$, the average size of the sawteeth in $(\ell_{N}^{t})_{t\geq 0}$ is significantly higher, in line with the steep increase of $\mathbb{E}\left[L_{N}\right]$ with $s$, see Figure 4. More importantly, the dynamics of $(c_{N}(t))_{t\geq 0}$ does not mirror that of $(\ell_{N}^{t})_{t\geq 0}$ any more. More precisely, a collapse of $(\ell_{N}^{t})_{t\geq 0}$ still implies a collapse of $(c_{N}(t))_{t\geq 0}$, simply because a tiny ancestral tree provides no opportunity to accumulate innovations, no matter how they are transmitted. However, $(c_{N}(t))_{t\geq 0}$ remains below $(\mu\ell_{N}^{t})_{t\geq 0}$, because inheritance of traits is increasingly diluted with their age.

### v2

@@ 3.5 Understanding the sawtooth: ancestral sets and descendant process
To understand the sawtooth behaviour observed in the supercritical case, that is, the linear increase of $(L_{N}^{t})_{t\geq 0}$ interrupted by sudden near-extinctions at random times, we need the finer picture of $\lvert\Lambda^{t}_{[n]}(\tau)\rvert_{\tau\geq 0}$ in Figure 14. The figure shows realisations of $\lvert\Lambda^{t_{i}}_{[N]}(t_{i}-t)\rvert_{t\in[t_{i}]_{0}}$ and $\lvert\Lambda^{t_{i}}_{\alpha}(t_{i}-t)\rvert_{t\in[t_{i}]_{0}}$ for $\alpha\in[N]$, $i\in\{1,2,3,4\}$, and starting times $t_{i}$ chosen so that
@@
where $t_{i}^{\rm e}$ is the time point where $(\Lambda^{t_{i}}_{[N]}(t_{i}-t))_{t\in[t_{i}]_{0}}$ goes extinct; so $t_{i}-t_{i}^{\rm e}$ is the corresponding realisation of $T^{t_{i}}$. Note that $T^{t_{2}},T^{t_{3}}$, and $T^{t_{4}}$ are coupled. With $\lesssim$ and $\ll$, we indicate that the quantities are close and not close to each other, respectively. We now see that the realisations of $(\Lambda^{t_{3}}_{[N]}(t_{3}-t))_{t\in[t_{3}]_{0}}$ and $(\Lambda^{t_{2}}_{[N]}(t_{2}-t))_{t\in[t_{2}]_{0}}$ quickly get close to, and actually become identical with, the realisation of $(\Lambda^{t_{4}}_{[N]}(t_{4}-t))_{t\in[t_{4}]_{0}}$; in particular, they join into the same metastable set and are extinguished with it. Likewise, the realisations of $(\Lambda^{t_{i}}_{\alpha}(t_{i}-t))_{t\in[t_{i}]_{0}}$, $i\in\{2,3,4\}$ and $\alpha\in[N]$, either die out quickly or merge with the metastable set. This merging is clear by arguments similar to those in Section 3.3: since $\Lambda^{t_{4}}_{[N]}(t_{4}-t_{3})\subseteq\Lambda^{t_{3}}_{[N]}(0)=[N]$, we have $\Lambda^{t_{4}}_{[N]}(t_{4}-t_{3}+\tau)\subseteq\Lambda^{t_{3}}_{[N]}(\tau)$ for all $\tau>0$; so, by moving towards the metastable size, the latter process joins into the former from above. Likewise, $\Lambda^{t_{3}}_{\alpha}(0)\subseteq\Lambda^{t_{3}}_{[N]}(0)$, so $\Lambda^{t_{3}}_{\alpha}(\tau)\subseteq\Lambda^{t_{3}}_{[N]}(\tau)$ for $\tau>0$; hence the former process, if it does not die out quickly, joins into the latter from below. Analogous arguments hold for the other time points.
@@
So far, we have followed the ancestral process backward in time. But the figure can also be read forward. Let $\Gamma_{[n]}^{t}(\varrho)$ be the set of descendants at time $t+\varrho$, $\varrho\geq 0$, of the set $[n]$ of individuals at forward time $t$. The process $(\Gamma_{[n]}^{t}(\varrho))_{\varrho\geq 0}$ has the same law as $(\Lambda_{[n]}^{t}(\tau))_{\tau\geq 0}$ with $\tau$ replaced by $\varrho$; this follows immediately from the fact that the set of descendants (ancestors) is obtained by following all learning arrows in the forward (backward) direction and pruning lines that meet a cross, and a reversal of all arrows does not change the law of the process. As a side remark, let us note that $\lvert\Gamma_{[n]}^{0}(t)\rvert_{t\geq 0}$ has the same law as $(X(t))_{t\geq 0}$ of our single-trait model with $X(0)=n$; this reflects the self duality of the SIS model. In any case, like $\lvert\Lambda_{[N]}^{t}(\tau)\rvert_{\tau\geq 0}$, also $\lvert\Gamma_{[N]}^{t}(\varrho)\rvert_{\varrho\geq 0}$ moves to the metastable state around $N\bar{\xi}$ quickly, and like $\lvert\Lambda_{\alpha}^{t}(\tau)\rvert_{\tau\geq 0}$, also $\lvert\Gamma_{\alpha}^{t}(\varrho)\rvert_{\varrho\geq 0}$ dies out quickly with probability $u/s$ and otherwise moves to the metastable state.
@@
$\alpha\in\Lambda_{\beta}^{t}(t-r)\Longleftrightarrow\beta\in\Gamma_{\alpha}^{r}(t-r),\quad\alpha,\beta\in[N],\;0\leq r\leq t,$
@@
$\Lambda_{[N]}^{t}(t-r)=\{\alpha:\Gamma_{\alpha}^{r}(t-r)\neq\varnothing\}.$ (32)
@@
In general, $\Lambda_{[N]}^{t}(t-r)\neq\Gamma_{[N]}^{r}(t-r)$ realisationwise (although they are equal in distribution). However, it is true that
$\Lambda_{[N]}^{t}(t-r)=\varnothing\Longleftrightarrow\Gamma_{[N]}^{r}(t-r)=\varnothing$ (33)
@@
for any $0<r<t$, because if the population at time $t$ does not have ancestors at time $r$, then the descendants of the population at time $r$ do not survive until $t$. Moreover, if $t\gg r$ and the sets in (33) are not empty, we have
$\lvert\Lambda_{[N]}^{t}(t-r)\rvert\approx N\bar{\xi}\approx\lvert\Gamma_{[N]}^{r}(t-r)\rvert$ (34)
@@
realisationwise due to metastability. In the realisation of Figure 14, the equalities in (33) are true for $r=t_{1}$, $t=t_{2}$; that is, both processes go extinct in $t_{1},t_{2}]$. In contrast, for $r=t_{2}$ and $t=t_{3}$ or $t=t_{4}$, (34) applies.
Now, (33) and (34) together allow us to approximate $L_{N}^{t}$ of (10) as
$L_{N}^{t}\approx\int_{-\infty}^{t}\lvert\Gamma_{[N]}^{r}(t-r)\rvert\,\mathrm{d}r=\int_{0}^{\infty}\lvert\Gamma_{[N]}^{t-\varrho}(\varrho)\rvert\,\mathrm{d}\varrho=\int_{0}^{S^{t}}\lvert\Gamma_{[N]}^{t-\varrho}(\varrho)\rvert\,\mathrm{d}\varrho\approx N\bar{\xi}S^{t},$ (35)
@@
where $S^{t}$ denotes the largest $\varrho\leq t$ such that $\Gamma_{[N]}^{t-\varrho}(\varrho)\neq\varnothing$, or, equivalently, the smallest $\varrho\leq t$ such that $\Gamma_{[N]}^{(t-\varrho)-}(\varrho)=\varnothing$. This is the time [^6] Note that $(\Gamma_{[n]}^{t}(\varrho))_{\varrho\geq 0}$ is based on the forward process, so it is càdlàg; that is, a jump at time $t$ means that the ‘old’ state applies until time $t$- (the moment ‘just before’ $t$), and at time $t$, the process is already in the ‘new’ state. This implies that the ancestors are extinct at $(t-S^{t})$-, but alive at $t-S^{t}$. since the last extinction event before $t$. In other words, the population at forward time $t$ has their oldest ancestor(s) at forward time $t-S^{t}$. Since the extinction times are distinct random points on the time axis, $(S^{t})_{t\geq 0}$ is a sawtooth function that increases linearly with slope 1 and is reset to some small value at random times. So (35) explains the linear increase in forward time, interrupted by collapses at the jumps of $(S^{t})_{t\geq 0}$.
Since $(\Gamma_{[N]}^{t}(\varrho))_{\varrho\geq 0}$ has the same law as $(\Lambda_{[N]}^{t}(\tau))_{\tau\geq 0}$, it is clear that $S^{t}$ has the same law as $T^{t}$, namely the exponential distribution with expectation $\mathbb{E}[T]$ of (29). Eq. (35) therefore implies that $\mathbb{E}[L_{N}]\approx N\bar{\xi}\mathbb{E}[T]$. As a consistency check, we compared this numerically with the exact value of $\mathbb{E}[L_{N}]$ of (23) for our parameter values $u=1,N=100$, and $s\in[0,1.4]$ and found good agreement (not shown).
@@
We can now also better understand the dynamics of $(L_{1}^{t})_{t\geq 0}$ and $(C_{1}(t))_{t\geq 0}$ in the lower panel of Figures 12 and 13. Since, for any given $\alpha$ and $t$, $(\Lambda^{t}_{\alpha}(\tau))_{\tau\geq 0}$ quickly either dies out or gets close to $(\Lambda^{t}_{N}(\tau))_{\tau\geq 0}$, it is clear that, most of the time, $(L_{1}^{t})_{t\geq 0}$ is either close to 0 or close to $(L_{N}^{t})_{t\geq 0}$. The frequent, random transitions between the two possibilities come from the fact that, at the metastable state over time, the actual set of lines that has old ancestors and thus ample knowledge, while being approximately constant in size, moves around in the population via the events of teaching individuals outside the current set, and by death events, thus rapidly including or excluding individuals. That is, by learning from someone with old ancestors, an individual acquires the entire ancestry and knowledge of the parent; and when the individual dies, it loses all its ancestry, including all its traits.
@@
Let us note in passing that, for the parameter values in Figure 12, (29) gives $\mathbb{E}\left[T\right]\approx 85.3$. Manual counting in Figure 12 yields $\approx 83$ sawteeth, which amounts to a mean tooth length (or extinction time) of 10000/83 $\approx$ 120. This overestimation of the length is presumably due to the underestimation of the number of teeth because small teeth are not resolved.
3.6 Related concepts and phenomena in population genetics
While the evolving genealogies considered here, and, in particular, their sawtooth dynamics, seem to be new, there are some related concepts and phenomena in population genetics. We have already mentioned the model of Muller’s ratchet with tournament selection (González Casanova et al., 2023), where evolving ASGs play a crucial role. In contrast to the ALG, these ASGs also have coalescence events, but the general behaviour is similar: on a longer time scale, there are repeated extinction events of the (so-called minimal-load) ASG (they correspond to the clicks of the ratchet). Between these events, the line-counting process of the ASG spends most of its time in a quasistationary state. On a shorter time scale between the extinction events, ASGs corresponding to different samples merge quickly, provided these samples are sufficiently large.
The connection to the aforementioned evolving genealogies in Kingman’s coalescent (Pfaffelhuber and Wakolbinger, 2006; Pfaffelhuber et al., 2011) is somewhat more remote. In contrast to the ALG, these genealogies are trees; their lengths also display a sawtooth behaviour (see Pfaffelhuber et al. (2011, Fig. 2)), but this stems from the change of roots (that is, the most recent common ancestor) of the trees (and subtrees), as opposed to the extinction of the ALG. Interestingly, Pfaffelhuber and Wakolbinger (2006, Sec. 5) also superpose an innovation mechanism, namely infinite-sites mutation, to their genealogies, which is perfectly analogous to our invention of traits. The mutations along the tree then translate into the differences between contemporary sequences, and the jumps of the most recent common ancestor lead to clusters of such sequence differences, in line with a classical observation by Watterson (1982).
@@ 3.7 Popularity spectrum
Aoki (2018) (see also Strimling et al. (2009); Fogarty et al. (2015, 2017)) studied the popularity spectrum $({P_{l}})_{l\in[N]}$ in discrete-time models of cultural evolution under various assumptions on the innovation and transmission mechanisms, where $P_{l}\mathrel{\mathop{:}}=\mathbb{E}\left[C_{N,l}\right]$ and $C_{n,l}$ is the number of traits carried by exactly $l$ individuals in a sample of size $n$ at stationarity; this was obtained via the equilibrium condition in the forward model. In Appendix D, we adapt this method to our model and derive
@@
We would like to note that a bimodal popularity spectrum is not tied to joint transmission of traits ($b=1$). It also occurs in the model of Nakamura et al. (2020) with independent transmission ($b<1$) when an individual may learn both from its genetic parent (vertical transmission) and from any other individual (horizontal transmission). Bimodality is, therefore, more than just an artifact of $b=1$.
@@ 4 Discussion
Mathematical models for the accumulation of cultural traits in a population over time are crucial for understanding the cultural variation in human populations. Here we have investigated cultural dynamics in which selectively neutral, discrete cultural traits are invented by individuals independently of one another at a constant rate and transmitted between individuals through random social learning. Related previous analyses (Strimling et al., 2009; Lehmann et al., 2011; Fogarty et al., 2015, 2017; Aoki, 2018; Nakamura et al., 2020) were mostly static and centered around the traits, focussing on the expectation of quantities like the number of distinct cultural traits maintained in a population at stationarity. Our approach is different in three ways: first, we concentrate on the underlying genealogies (as first described by Aguilar and Ghirlanda (2015)) and thus obtain results that remain true independently of the traits. Second, we investigate the dynamical aspects, which seem to be unexplored so far. At the heart of this are the time evolution of the set of ancestors of a single individual or the entire population; the merging process between the various ancestral sets; the metastability of the corresponding counting processes; and the concept of evolving genealogies. Third, we also obtain moments of $L_{n}$ and $C_{n}$.
@@
To keep the mean number of traits transmitted via learning events constant, we replace $s$ by $s/b$. For $b=0.99$ and, hence, our usual choice $s=1.3$ replaced by $s=130/99$, $(c_{N}(t))_{t\geq 0}$ still follows $\mu(\ell_{N}^{t})_{t\geq 0}$ closely, see Figure 16 (top). For $b=13/14$ and, hence, $s=1.3$ replaced by $s=1.4$ (Figure 16 bottom), the average size of the sawteeth in $(\ell_{N}^{t})_{t\geq 0}$ is significantly higher, in line with the steep increase of $\mathbb{E}\left[L_{N}\right]$ with $s$, see Figure 4. More importantly, the dynamics of $(c_{N}(t))_{t\geq 0}$ does not mirror that of $(\ell_{N}^{t})_{t\geq 0}$ any more. More precisely, a collapse of $(\ell_{N}^{t})_{t\geq 0}$ still implies a collapse of $(c_{N}(t))_{t\geq 0}$, simply because an extremely short genealogy provides no opportunity to accumulate innovations, no matter how they are transmitted. However, $(c_{N}(t))_{t\geq 0}$ remains below $(\mu\ell_{N}^{t})_{t\geq 0}$, because inheritance of traits is increasingly diluted with their age.
@@
The aim of the present study was to develop the genealogical theory of cultural evolution, highlighting its unique features as well as close relationship with ancestral structures in population genetics. The cultural genealogies form the basis upon which arbitrary trait transmission models may be superposed; ours was chosen for its mathematical appeal rather than for realism or even data analysis.
For the more realistic transmission model mentioned above (with independent transmission both horizontal and vertical), Nakamura et al. (2020) have compared the predicted popularity spectrum to the empirical equivalent extracted from datasets of cultural variation between archaeological sites or ethno-linguistic groups. Such datasets record the variation in, for example, bead types in Palaeolithic Europe, pottery in Neolithic Europe, and Fijian and Polynesian canoe design in binary format (presence and absence of each character) at the level of groups of humans rather than single ones. (The number of groups in these data sets ranges from 10 to 195. Incidentally, therefore, $N=100$ as chosen for computational practicality in our simulations seems to be appropriate.) A note of caution is in order, however: if there is an inherent temporal dynamics, as in our model, then the pure equilibrium consideration and use of cross-sectional data alone may cause misleading results. However, the temporal resolution required for a comparison with the dynamics of learning models cannot be expected in archeological or ethnographic data. Possibly, digital media (e.g., social network services) may provide such data for future work.